Circular No. 5882

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET URATA-NIIJIMA (1993q) This comet has been recovered by J. V. Scotti with the Spacewatch

telescope at Kitt Peak, as shown below. A 11′′ coma was visible on Oct. 20.49 (15′′ on Oct. 21.50); a 0′.4 tail was visible on both nights in p.a. 290^{o}- 292^{o}; m_{2} = 22.7 on both nights. The indicated correction to the prediction on MPC 16380 (ephemeris on MPC 22064) is ΔT = +0.24 day.

1993 UT  α_{2000}  δ_{2000}  m_{1}
Oct. 20.48992  10^{h}24^{m}24.^{s}33  +33^{o}03′28′′.7  19.4
20.49859  10 24 25.51  +33 03 26.5
20.50731  10 24 26.76  +33 03 23.1  19.2
21.48952  10 26 50.29  +32 58 20.4
21.49821  10 26 51.65  +32 58 17.0  19.6

NOVA AQUILAE 1993 C. E. Woodward and J. Cole, University of Wyoming; and R. D. Gehrz,

University of Minnesota, report the following infrared broadband magnitudes, obtained using the WIRO 2.3-m telescope (+ University of Minnesota bolometer and remote computer link) on Oct. 6.1 UT: J = 9.67 ± 0.44, H = 8.64 ± 0.06, K = 6.21 ± 0.04, L = 3.45 ± 0.01, M = 2.40 ± 0.01, N = 1.02 ± 0.07. Additional photometry through the narrow-band IRTF silicate filter set gave magnitudes of [7.8 μm] = 1.31 ± 0.06, [8.7 μm] = 1.23 ± 0.03, [9.8 μm] = 1.13 ± 0.04, [10.3 μm] = 1.06 ± 0.05, [11.6 μm] = 1.77 ± 0.16, and [12.5 μm] = 2.81 ± 0.17. These data are consistent with continuum emission from a 750-K blackbody and suggest that dust has formed in the ejecta (cf. IAUC 5831). There is no evidence for a strong [Ne II] emission line at 12.8 μm.

SUPERNOVA 1993ac IN ANONYMOUS GALAXY B. Schmidt and R. Kirshner, Harvard-Smithsonian Center for Astro-

physics (CfA), report that inspection of a spectrogram taken by J. Huchra (CfA) and J. Brodie (University of California, Santa Cruz) at the Multiple-Mirror Telescope on Oct. 20.5 UT reveals that this object is a type-Ia supernovae ∼ 1 week past maximum — not a type-II supernova as hinted on IAUC 5879. The redshift derived from the spectrum of the host galaxy is 14 800 km/s.

1993 October 21 Daniel W. E. Green